If the standard reduction potential for A --> A+ + e- was 1.21 volts and the standard reduction potential for B --> B+2 + 2 e- was 0.46 volts, what would be the standard potential for the following reaction: B+2 + 2 A --> 2 A+ + B?

If the standard reduction potential for X --> X+2 + 2 e- was 0.53 volts and the standard potential for the reaction: X + Y+2 --> X+2 + Y was 0.4 volts, what would be the standard reduction potential for: Y --> Y+2 + 2 e-?

What element is undergoing oxidation in the following reaction: Cr2O3 + Al --> Al2O3 + Cr

Which element is undergoing reduction in the following reaction: FeF2 + CrF3 --> CrF2 + FeF3? (Give the name.)
Answer #1

oues 2222222 a o g²+ + 2A t é Bettzé QAT+B Eu 1.21 v = 0.46 E = ? Solution e canderd o sincer electrode potential is intensivGiven that A at t é ti = 10210 Here = 7 Let Gibbs free energy og, = - nf El => ag, = -1 xfxfi multiplying by 2 in both side(het) Now B - B²+ tre o E = o.uGuia Reversing this equation : 7 B²+ trê BED E 8 = -0.46N 13=2 ст-Р (2) 2A 2Attre ® E 2 = 1021E aius x x² + trêt = 0.53v Oglannaf Ei & x+y +2+ + y = 0uur I y - y2+ tzł = ? Reversing equation number ) :- x2+ ty t y 2+ EO7 + 092 09 -) enfeit (haft - 26% -) - 2(0-53) + 12 F*(-0.44 ] = 426 € - 0.53+0.4 = -6° -) FO = +0.53 -0.4 - Answer • Ques ♡ +of Al moves from • But onudation state wio to to. Thus All is undergoing oncidation. - Answer +2 Fefa +2 +3 t + 3 ex fg - Cs